OpenStudy (anonymous):

Which line is parallel to the line that passes through the points (-1, 2) and (5, -4)?

OpenStudy (anonymous):

eq. of line is \[y-y _{1}=\frac{ y _{2}-y _{1} }{x _{2}-x _{1} }\left( x-x _{1} \right)\] solve it and write in the form ax+by=c Then line parallel to ax+by=c is ax+by=d ,where d is a constant.

OpenStudy (anonymous):

1/3